# A Langevin sampling algorithm inspired by the Adam optimizer

### v1

@@ 4 Adam-inspired Monitor Function and Sundman Transform
The $\zeta$-dynamics (9) then becomes
@@
An Adam-like choice of $g$ is $g(x,p)=\|\nabla U(x)\|^{2}$, i.e., monitoring the force norm. The closest fidelity to (22) is thus found with $r=1/4$ in the Sundman transformation, since scaling the full vector field by $a$ can be related to scaling half the system only by $a^{2}$. For flexibility we propose to use $g(x,p)=\Omega^{-1}\|\nabla U(x)\|^{s}$, i.e. monitoring the force norm raised to some possibly fractional, positive power $s$ and scaled by a normalization factor $\Omega^{-1}$. Effectively, this choice combined with raising $\zeta$ to the $r$th power in the restriction function is tantamount to controlling the timestep based on the $r\cdot s$ power of the gradient norm.
Where it is not a physical parameter of the modelling task, the temperature $\beta^{-1}$ allows the method to behave more like an optimizer ($\beta^{-1}\approx 0$) or a sampling scheme ($\beta^{-1}\rightarrow 0$). In a Bayesian sampling context, $U(x)$ is the negative log-posterior and we may take $\beta^{-1}=1$ to sample the posterior or $\beta^{-1}<1$ to implement annealed importance sampling [72]. Due to the various points mentioned above, SamAdams with $s\cdot r=1/2$ does not strictly reduce to Adam for $\beta^{-1}\rightarrow 0$. In our experiments we found that the optimal choices of $r$ and $s$, as well as other coefficients, were problem-class dependent, but within a specific class of models those selections were relatively easy to decide.
@@ 4.1 Other Monitor Functions
The choice to make SamAdams relate to the Adam optimizer is not unique: possible alternatives include basing $g$ on only the prior in a Bayesian sampling setting or basing $g$ on some subset of the variables (which are known to exhibit high levels of variability). It would also be possible to introduce higher derivative information derived from the potential energy function such as the trace or determinant of the Hessian matrix. We have not substantially explored such options. For noisy gradient evaluations, one might also design a monitor function based on the estimation of gradient noise such that stepsizes are decreased whenever the injected gradient noise is large. This idea is motivated in Appendix D, which shows results for logistic regression with stochastic gradients, during which the adaptive stepsize strongly reacts to the employed batch size.
@@ 5.1 Asymmetric Double Well
We first demonstrate the stepsize adaptation of SamAdams on a one-dimensional double well problem where one well is much narrower than the other. The potential is given by $U(x)=\frac{b}{L}(x+1)^{2}(x-L)^{6}$ with $b=1.5$ and $L=2$. We pick a small temperature $T\equiv\beta^{-1}=0.4$ to make the transition across the barrier reasonably rare. In the narrow well, a Langevin path will encounter greater forces, causing SamAdams to decrease its stepsize $\Delta t$ accordingly. Fig. 3 (top) shows the potential and the density on the left, and the $x$- and $\Delta t$-values of a single trajectory of SamAdams on the right. We clearly see that occupation in the narrow well (negative $x$-values) leads to a restriction of the stepsize $\Delta t$ to small values.
@@
Figure 3: Sampling experiments on a 1D toy model (see text). (a) Potential and density for employed temperature $T=0.4$. (b) $x$-coordinate and adaptive stepsize $\Delta t$ for SamAdams along a single trajectory. The black dashed line gives the value of virtual stepsize $\Delta\tau$ which is adaptively increased or reduced to yield the real stepsize $\Delta t$. (c) Mean absolute errors of two observables, the $x$-coordinate and the occupation frequency of area $x<0.5$ against (mean) stepsize $\Delta t$. The different values for SamAdams were obtained by varying $\Delta\tau$ from 0.03 to 0.2. (d) $\Delta t$ histograms for SamAdams run at three different $\Delta\tau$. Other hyperparameters: $\gamma=\alpha_{1}=\alpha_{2}=1$, $m=0.1$, $M=10$, $r=0.25$, $s=2$.
@@
The lower left panel of Fig. 3 shows absolute mean errors of two observables, the $x$-coordinate and indicator function of the domain $x<0.5$ (which roughly corresponds to the occupation probabilities of the barrier and the narrow well, respectively), against stepsize (mean stepsize for the adaptive scheme). We first ran SamAdams for different values $\Delta\tau$, measured the mean adaptive stepsizes $\langle\Delta t\rangle$ for each of these runs, and then ran BAOAB at stepsizes fixed to these values to obtain canonical averages at the same mean stepsize and compute cost. Each point in the figure was generated by averaging over 300 independent trajectories for $5\cdot 10^{7}$ iterations (discarding the first 100,000 iterations as burn-in). The ground truths $\langle x\rangle$ and $P(x<0.5)$ were obtained via numerical quadrature. We also plot vertical lines denoting the maximum (mean) stepsize at which the corresponding algorithm became unstable in at least one of the 300 trajectories (i.e., the stability threshold). We see that BAOAB does significantly worse than SamAdams, only reaching similar accuracies for the smallest stepsize examined. SamAdams’ performance barely depends on $\langle\Delta t\rangle$ until very close to its stability threshold. We also observe that it is able to use larger steps than BAOAB.
@@
The bottom right of Fig. 3 shows the $\Delta t$ histograms for three of the SamAdams runs. They have a bimodal structure, as might be expected from an asymmetric double well. All three of them are able to use stepsizes larger than the stability threshold of BAOAB, supporting the idea that the stability threshold for fixed-stepsize schemes depends on local variation of the loss landscape, rather than on the landscape as a whole. As for the star potential of Fig. 1, it is enough to use small stepsizes in critical areas. Note how one of the $\Delta t$ histograms has a mean of $\langle\Delta t\rangle=0.216$, which is 31% larger than BAOAB’s threshold, which together with the low observable errors even at that stepsize implies a substantial increase in computational efficiency compared to BAOAB. The ability to use larger stepsizes (in some cases much larger stepsizes) than constant stepsize schemes while preserving sampling quality will also be observed in the examples of the next subsection.
@@ 5.2.1 Neal’s Funnel
Our goal is to sample the canonical distribution at temperature $T=1$. We used $\gamma=5$ and discarded $10^{5}$ steps as burn-in (equilibration). The initial position was taken to be $(0,5)$ in a relatively flat zone, and we set initial momenta and $\zeta_{0}=0$.
@@
Figure 4: (left and center) A BAOAB trajectory with stepsize $\Delta t=0.04$ shows an unstable evolution in $10^{6}$ steps. The descent into the funnel leads to a spike in both kinetic temperature and mean potential energy. Although shortlived, this type of event can, as here, corrupt long term averages. (right) At longer times, these instabilities are inevitable, for stepsizes above or equal to $\Delta t=0.015$.
By contrast, SamAdams (ZBAOABZ) produces reliable, stable trajectories with much larger mean stepsize than any fixed stepsize method. In Fig. 5 we show a trajectory with mean stepsize $\langle\Delta t\rangle=0.16$ ($N=10^{7}$). (The details of stepsize variation are as follows: $\Delta t_{\rm min}=10^{-4}$, $\Delta t_{\rm max}=0.6$, $\Delta t_{0}=0.6$, $\alpha=0.1$, $r=0.5$ and $g(x,p)=\|\nabla U\|$. We used the second form of the filter function $\psi=\psi^{(2)}$ in all the 2d examples.)
A histogram computed using the samples obtained from SamAdams is virtually identical to the target distribution (Fig. 6), despite requiring around 10% of the computational effort needed if the corresponding fixed-stepsize method was used. Observables, e.g. the kinetic temperature or the mean potential energy as shown in Fig.5 are approximated to three significant digits.
@@
Figure 6: Here we compare the (weighted) histogram of solution data to the actual canonical distribution for the SamAdams trajectory with $\langle\Delta t\rangle=0.16$. The distributions are visually very similar.
@@ 5.2.2 Entropic Barrier
We simulate this with temperature set to $0.05$ to create a challenging test. We used friction $\gamma=5$ in this example which gave approximately optimal results.The initial point was taken at $(x,y)=(3,0)$ near the minimum on the right side, and the initial momenta were set to zero. The SamAdams parameters were like in the previous example ($r=0.5$, $\alpha=0.1$), but we set $\Delta t_{\rm min}=0.0001$, $\Delta t_{\rm max}=0.5$ $\Delta t_{0}=0.5$.
Figure 7: Comparison of trajectories obtained using fixed and variable stepsize. (a) BAOAB with fixed stepsize $\Delta t=0.1$ converges as expected, with around 2-3 barrier crossings per 1M steps. (b) BAOAB becomes unstable above $\Delta t=0.15$ and at $\Delta t=0.2$ shows no diffusion over the barrier. (c) A variable stepsize trajectory with $\langle\Delta t\rangle=0.356$, restoring the performance of small fixed stepsize.
Figure 8: Large variable stepsize results ($\langle\Delta t\rangle=0.356$). In the left figure, a histogram of the trajectory data (weighted by the time-rescaling used to generate the data) is shown. This can be compared to the exact distribution (from the probability density function) shown in the central panel. Finally the actual stepsize distribution is given in the right panel and shows that a vanishingly small number of steps require a small (below 0.2) stepsize. Not visible in the histogram is the fact that a very small number of steps used a stepsize below 0.01.
@@
In the left panel of Fig. 8 we show a (weighted) histogram of the computed states for $\langle\Delta t\rangle=0.356$ which can be compared with the exact distribution in the central panel. In this example, with $10^{8}$ steps, the kinetic and configurational temperatures (observables which should each average to the target temperature [59]) are accurate to within 1% ($T_{\rm kin}=0.04947$, $T_{\rm conf}=0.04954$). Of particular interest is the fact that the stepsize distribution shown at right in Fig. 8 has barely any mass below $\Delta t=0.2$, meaning that the small stepsizes are only needed at very rare instances of barrier crossing (precisely in a small neighborhood of the origin).
@@ 5.2.3 Beale Potential
We used a temperature of $T=3$ to obtain sufficient barrier transitions. The initial point was $(x,y)=(3,0)$, with zero momenta.
@@
For fixed stepsize integration, the step needs to be below about 0.003. Running at or above this threshold almost all runs of $10^{8}$ steps result in failure. At the small value $\Delta t=0.0025$, $10^{6}$ samples are not enough to cover the distribution, as shown in Fig. 9. By contrast (see the right panel of Fig. 9) we were able to use SamAdams (ZBAOABZ) reliably with a mean step of $\Delta t=0.022$ which indicates a stability improvement of nearly an order of magnitude. While there is some visible error in the histogram for SamAdams (Fig. 9) it is much less than for BAOAB operating at its much smaller stable stepsize.
@@
Figure 9: For the Beale potential, we show the probability distribution at the left and the histogram (center) for a fixed stepsize BAOAB run with $\Delta t=0.0025$ (larger stepsizes are unreliable), $N=10^{6}$ steps. Finally, at right we see a histogram obtained from a SamAdams trajectory $\langle\Delta t\rangle=0.022$, $N=10^{6}$ steps. Parameters: $\Delta t_{\rm min}=0.001$, $\Delta t_{\rm max}=0.1$, $\Delta t_{0}=0.1$ $r=0.5$, $\alpha=1$, and monitor function $g(x,p)=\|\nabla U(x)\|$.
@@
We also perform more extensive runs to resolve the bias in canonical averages. The procedure is similar to the case of the 1D toymodel in Sec. 5.1. To obtain different $\langle\Delta t\rangle$ for SamAdams, we varied $\Delta\tau$, keeping everything else fixed, and then ran BAOAB at fixed stepsizes set to the obtained $\langle\Delta t\rangle$ values for comparison. We averaged over 200 independent trajectories, where the number of iterations scaled with $\Delta\tau$ (we used $10^{8}$ iterations at $\Delta\tau=0.001$, and varied $\Delta\tau$ from 0.00025 to 0.06). The computational cost of each BAOAB simulation was the same as for the corresponding SamAdams run. Fig. 10 shows the bias against (mean) stepsize $\Delta t$ for the two temperatures and coordinates (the ground truths for the latter were obtained via numerical quadrature). While SamAdams and BAOAB show comparable errors for small stepsizes, BAOAB becomes unstable at $\Delta t=0.0025$. SamAdams can still be run at $\langle\Delta t\rangle=0.01$ without any decline in accuracy. If a modest decline in accuracy is acceptable, it can be run at substantially larger steps. In fact, while the BAOAB curves in Fig. 10 stop at the last stable stepsize, SamAdams remained stable until the last examined $\Delta\tau$.

### v2

@@ 4 Adam-inspired Monitor Function and Sundman Transform
Remark 3.
The choice 23 for the function $g$ is not globally Lipschitz, and does not satisfy the assumptions of Theorem 1 because of the square. However, one could also consider $g$ to be the norm of $\nabla U$ with power $r$ (which is then globally Lipschitz for $r\leq 1$, but has a singularity at points where $\nabla U$ vanishes). To fulfil the assumptions, you could also use instead the smoothed version $(\|\nabla U\|^{2}+\epsilon)^{s/2}$, $s\leq 1$.
In practice, we could instead consider $\psi(\zeta)$ rather than $\zeta$ as a variable in the reweighting and the SDE, which is uniformly bounded from above and below. The time-rescaled process (10)-(13) could be expressed in this way.
With (23) the $\zeta$-dynamics (9) becomes
@@
An Adam-like choice of $g$ is $g(x,p)=\|\nabla U(x)\|^{2}$, i.e., monitoring the force norm. The closest fidelity to (22) is thus found with $r=1/4$ in the Sundman transformation, since scaling the full vector field by $a$ can be related to scaling half the system only by $a^{2}$. For flexibility, we propose to use
$g(x,p)=\Omega^{-1}\|\nabla U(x)\|^{s},$ (26)
monitoring the force norm raised to some possibly fractional, positive power $s$ and scaled by a normalization factor $\Omega^{-1}$. Effectively, this choice combined with raising $\zeta$ to the $r$th power in the restriction function is tantamount to controlling the timestep based on the $r\cdot s$ power of the gradient norm.
Where it is not a physical parameter of the modelling task, the temperature $\beta^{-1}$ allows the method to behave more like an optimizer ($\beta^{-1}\approx 0$) or a sampling scheme ($\beta^{-1}>0$). In a Bayesian sampling context, $U(x)$ is the negative log-posterior and we may take $\beta^{-1}=1$ to sample the posterior or $\beta^{-1}<1$ to implement annealed importance sampling [74]. Due to the various points mentioned above, SamAdams with $s\cdot r=1/2$ does not strictly reduce to Adam for $\beta^{-1}\rightarrow 0$. In our experiments we found that the optimal choices of $r$ and $s$, as well as other coefficients, were problem-class dependent, but within a specific class of models those selections were relatively easy to decide.
@@ 4.1 Other Monitor Functions
The choice to make SamAdams relate to the Adam optimizer is not unique: possible alternatives include basing $g$ on only the prior in a Bayesian sampling setting or basing $g$ on some subset of the variables (which are known to exhibit high levels of variability). It would also be possible to introduce higher derivative information derived from the potential energy function such as the trace or determinant of the Hessian matrix. We have not substantially explored such options. For noisy gradient evaluations, one might also design a monitor function based on the estimation of gradient noise such that stepsizes are decreased whenever the injected gradient noise is large. This idea is motivated in Appendix F, which shows results for logistic regression with stochastic gradients, during which the adaptive stepsize strongly reacts to the employed batch size.
@@ 5 Numerical Experiments
In all our numerical experiments, we use the monitor function (26) for different $\Omega$ and $s$ (mostly $s\in\{1,2\}$), and the Sundman kernel $\psi$ may either be $\psi^{(1)}$ or $\psi^{(2)}$ from (25), with different values for $m$, $M$, and $r$. We note that the detailed finetuning of these hyperparameters is usually not necessary. As discussed in Sec. 3.1, there are various ways in which $\zeta_{0}$ may be initialized, and we make use of both presented variants. We examine the behavior of SamAdams in terms of individual trajectories as well as ergodic averages. In the case of the latter, we typically run multiple independent trajectories and first compute the intra-trajectory averages (i.e., time averages using the reweighting from Sec. 2.1) and then the inter-trajectory average. To measure sampling biases, we compute weak errors of observables $\phi(q)$, $q:=(x,p)$, defined by
$\lim_{n\to\infty}\Big{|}\mathbb{E}\big{[}\phi\big{(}q_{n}\big{)}\big{]}-\mathbb{E}\big{[}\phi\big{(}q(n\Delta t)\big{)}\big{]}\Big{|},$ (27)
where $q_{n}$ gives the iterates obtained by the simulation. The expectation of the numerical iterates, $\mathbb{E}\big{[}\phi\big{(}q_{n}\big{)}\big{]}$, is estimated by time- and trajectory averages obtained from simulation, where we sometimes use the notation $\langle\phi\rangle$ to denote that empirical average. We also write $\langle\Delta t\rangle$ to denote the empirical mean of the adaptive stepsize used by SamAdams in a given experiment. We make use of the two temperature observables often monitored in molecular dynamics[60]:
$T_{\text{kin}}$ $:=\frac{1}{N_{\text{dof}}}\|p\|^{2},$ (kinetic temperature) (28)
$T_{\text{conf}}$ $:=\frac{1}{N_{\text{dof}}}x\cdot\nabla U(x),$ (configurational temperature) (29)
with $N_{\text{dof}}$ as the number of degrees of freedom. The expectation of these quantities is known and given by the system temperature, $\mathbb{E}(T_{kin})=\mathbb{E}(T_{conf})=T=\beta^{-1}$. We also use other observables in our low-dimensional experiments whose expectations can be computed via numerical quadrature. Thus, the ground truths $\lim_{n\to\infty}\mathbb{E}[\phi(q(n\Delta t))]$ of the observable means are either known or easily obtainable. For the Bayesian neural network problems we do not have access to the true posterior means, but we measure sampling quality in terms of negative loss likelihood (given by the loss function) and posterior-averaged classification accuracies.
[…]
As mentioned in Sec. 3.1, there are various ways to initialize $\zeta$. is by simply setting $\zeta_{0}=0$. Since $\psi(\zeta_{0})=M$ for either of the given choices of the filter function, we have $\Delta t_{0}=M\Delta\tau$=$\Delta t_{\max}$. The advantage of this approach is that the initial stepsize is known, independently of the value of the driving function $g(x_{0},p_{0})$. The disadvantage is that the stepsizes $\Delta t_{n}$ can then never become larger than $\Delta t_{0}$, which may prevent efficiency gains in settings where the simulation is started in a critical area, requiring small stepsizes, and then moves into more benign areas, allowing for larger stepsizes. This approach is particularly suitable for cases in which one can pick the initial configuration $(x_{0},p_{0})$ such that $g(x_{0},p_{0})$ is small. In low-dimensional models this is often easy, and we have followed this procedure below. For higher-dimensional models one may first need to run an optimization scheme. Another way of initializing $\zeta$ is by setting $\zeta_{0}=g(x_{0},p_{0})$. That way, the initial stepsize $\Delta t_{0}$ is fully determined by the driving function at the initial conditions, which is natural given that $\Delta t_{n}$ with $n>0$ will depend on the current and past values of the driving function. This allows the algorithm to pick the “right” initial stepsize automatically, which will typically not lead to $\Delta t_{0}=\Delta t_{\max}$. The downside of this approach is that the user has no way to enforce a certain initial stepsize. It’s all up to the algorithm to pick one in the interval $(m\Delta\tau,\ M\Delta\tau]$.
The experiments compare ZBAOABZ to the constant-stepsize equivalent BAOAB. Since BAOAB is a high-quality sampling method in and of itself, able to outperform many of the conventionally used unadjusted schemes[56], showing that ZBAOABZ is able to outperform BAOAB is equivalent to pushing the limits of state-of-the-art sampling.
@@ 5.1 Asymmetric Double Well
We first demonstrate the stepsize adaptation of SamAdams on a one-dimensional double well problem where one well is much narrower than the other. The potential is given by $U(x)=\frac{b}{L}(x+1)^{2}(x-L)^{6}$ with $b=1.5$ and $L=2$. We pick a small temperature $T\equiv\beta^{-1}=0.4$ to make the transition across the barrier reasonably rare. Fig. 3a) shows the potential and the density. The hyperparameters of monitoring function (26) are $s=2$ and $\Omega=1$. We initialize $(x_{0},p_{0})=(2,0)$ and SamAdams variable $\zeta_{0}=g(x_{0},p_{0})=\|\nabla U(x_{0})\|^{2}$. We use Sundman kernel $\psi=\psi^{(1)}$ with $m=0.1$, $M=10$, $r=0.25$. Langevin friction and $\zeta$ friction are given by $\gamma=\alpha=1$. Since a Langevin path will encounter greater forces in the narrower potential well, the aim is for SamAdams to decrease its stepsize $\Delta t$ accordingly. That this is indeed the case can be seen in Fig. 3b), which shows the $x$- and $\Delta t$-values of a single SamAdams trajectory. It can be seen that occupation in the narrow well (negative $x$-values) leads to a restriction of the stepsize $\Delta t$ to small values.
@@
Figure 3: Sampling experiments on a 1D toy model (see text). a) Potential and Gibbs density for employed temperature $T=0.4$. b) $x$-coordinate and adaptive stepsize $\Delta t$ for SamAdams along a single trajectory. The black dashed line gives the value of virtual stepsize $\Delta\tau$ which is adaptively increased or reduced to yield the real stepsize $\Delta t$. c) Absolute mean errors of two observables, the $x$-coordinate and the occupation frequency of area $x<0.5$ against (mean) stepsize $\Delta t$. The different values for SamAdams were obtained by varying $\Delta\tau$ from 0.03 to 0.2. d) $\Delta t$ histograms for SamAdams run at three different $\Delta\tau$.
@@
Fig. 3c) shows weak errors of two observables, the $x$-coordinate and indicator function of the domain $x<0.5$ (which roughly corresponds to the occupation probabilities of the narrow well and the barrier), against stepsize (mean stepsize for the adaptive scheme). We first ran SamAdams for different values $\Delta\tau$, measured the mean adaptive stepsizes $\langle\Delta t\rangle$ for each of these runs, and then ran BAOAB at stepsizes fixed to these values to obtain canonical averages at the same mean stepsize and compute cost. Each point in the figure was generated by averaging over 300 independent trajectories for $5\cdot 10^{7}$ iterations (discarding the first 100,000 iterations as burn-in). We also plot vertical lines denoting the maximum (mean) stepsize at which the corresponding algorithm became unstable in at least one of the 300 trajectories (i.e., the stability threshold). We see that BAOAB does significantly worse than SamAdams, only reaching similar accuracies for the smallest stepsize examined. SamAdams’ performance barely depends on $\langle\Delta t\rangle$ until very close to its stability threshold. We also observe that it is able to use larger steps than BAOAB.
@@
Fig. 3d) shows the $\Delta t$ histograms for three of the SamAdams runs. They have a bimodal structure, as might be expected from an asymmetric double well. All three of them are able to use stepsizes larger than the stability threshold of BAOAB, supporting the idea that the stability threshold for fixed-stepsize schemes depends on local variation of the loss landscape, rather than on the landscape as a whole. As for the star potential of Fig. 1, it is enough to use small stepsizes in critical areas. Note how one of the $\Delta t$ histograms has a mean of $\langle\Delta t\rangle=0.216$, which is 31% larger than BAOAB’s threshold, which together with the low observable errors even at that stepsize implies a substantial increase in computational efficiency compared to BAOAB. The ability to use larger stepsizes (in some cases much larger stepsizes) than constant stepsize schemes while preserving sampling quality will also be observed in the examples of the next subsection.
@@ 5.2.1 Neal’s Funnel
Our goal is to sample the canonical distribution at temperature $T=1$. We used $\gamma=5$ and discarded $10^{5}$ steps as burn-in (equilibration). The initial position was taken to be $(0,5)$ in a relatively flat zone, and we set initial momenta and $\zeta_{0}$ to zero.
@@
Figure 4: Left, Center: A BAOAB trajectory with stepsize $\Delta t=0.04$ shows an unstable evolution in $10^{6}$ steps. The descent into the funnel leads to a spike in both kinetic temperature and mean potential energy. Although shortlived, this type of event can, as here, corrupt long term averages. Right: At longer times, these instabilities are inevitable, for stepsizes above or equal to $\Delta t=0.015$.
By contrast, SamAdams (ZBAOABZ) produces reliable, stable trajectories with much larger mean stepsize than any fixed stepsize method. In Fig. 5 we show a trajectory with mean stepsize $\langle\Delta t\rangle=0.16$ ($N=10^{7}$). (The details of stepsize variation are as follows: $\Delta t_{\rm min}=m\Delta_{\tau}=10^{-4}$, $\Delta t_{\rm max}=M\Delta_{\tau}=\Delta t_{0}=0.6$, $\zeta_{0}=0$, $\alpha=0.1$, $r=0.5$ and $g(x,p)=\|\nabla U\|$. We used the second form of the filter function $\psi=\psi^{(2)}$.)
A histogram computed using the samples obtained from SamAdams is virtually identical to the target distribution (Fig. 6), despite requiring around 10% of the computational effort needed if the corresponding fixed-stepsize method was used. Note that since a histogram can be interpreted as an observable itself, the samples obtained via SamAdams have to be reweighted as described in Sec. 2.1 before plotting the histogram. Observables like the kinetic temperature or the mean potential energy as shown in Fig.5 are approximated to three significant digits.
@@
Figure 6: Here we compare the (reweighted, see text) histogram of trajectory data for SamAdams to the actual canonical distribution. The distributions are visually very similar.
@@ 5.2.2 Entropic Barrier
We simulate this with temperature set to $0.05$ to create a challenging test. We used friction $\gamma=5$ in this example which gave approximately optimal results. The initial point was taken at $(x,y)=(3,0)$ near the minimum on the right side, and the initial momenta were set to zero. The SamAdams parameters were like in the previous example, but we set $\Delta t_{\rm min}=0.0001$, $\Delta t_{\rm max}=\Delta t_{0}=0.5$.
Figure 7: Comparison of trajectories obtained using fixed and variable stepsize. Left: BAOAB with fixed stepsize $\Delta t=0.1$ converges as expected (with around 2-3 barrier crossings per 1M steps). Center: BAOAB becomes unstable above $\Delta t=0.15$ and at $\Delta t=0.2$ shows no diffusion over the barrier. Right: A variable stepsize trajectory with $\langle\Delta t\rangle=0.356$, restoring the performance of small fixed stepsize.
Figure 8: SamAdams trajectory with large mean stepsize ($\langle\Delta t\rangle=0.356$). Left: Histogram of trajectory data. Center: Exact distribution. Right: Distribution of adaptive stepsize $\Delta t$. Only a vanishingly small number of steps require a small (below 0.2) stepsize. Not visible in the histogram is the fact that a very small number of steps used a stepsize below 0.01.
@@
In the left panel of Fig. 8 we show a histogram of the computed states for $\langle\Delta t\rangle=0.356$ which can be compared with the exact distribution in the central panel. In this example, with $10^{8}$ steps, the kinetic and configurational temperatures are accurate to within 1% ($T_{\rm kin}=0.04947$, $T_{\rm conf}=0.04954$). Of particular interest is the fact that the stepsize distribution shown at right in Fig. 8 has barely any mass below $\Delta t=0.2$, meaning that the small stepsizes are only needed at very rare instances of barrier crossing (precisely in a small neighborhood of the origin).
@@ 5.2.3 Beale Potential
We used a temperature of $T=3$ to obtain sufficient barrier transitions. The initial point was $(x,y)=(3,0)$, with zero momenta, and we use $\gamma=1$.
@@
For fixed stepsize integration, the step needs to be below about 0.003. Running at or above this threshold almost all runs of $10^{8}$ steps result in failure. At the small value $\Delta t=0.0025$, $10^{6}$ steps are not enough to cover the distribution, as shown in the center panel of Fig. 9. By contrast (see the right panel of Fig. 9) we were able to use SamAdams reliably with a mean step of $\langle\Delta t\rangle=0.022$ which indicates a stability improvement of nearly an order of magnitude. While there is some visible error in the histogram for SamAdams (Fig. 9) it is much less than for BAOAB operating at its much smaller stable stepsize.
@@
Figure 9: Sampling results for the Beale potential. Left: Gibbs distribution (ground truth). Center: Histogram for a fixed stepsize BAOAB run with $\Delta t=0.0025$ (larger stepsizes are unreliable). Right: Histogram obtained from a SamAdams trajectory with $\langle\Delta t\rangle=0.022$. Settings: $\psi=\psi^{(2)}$ with $\Delta t_{\rm min}=m\Delta\tau=0.001$, $\Delta t_{\rm max}=M\Delta\tau=0.1$, $\Delta t_{0}=0.1$, $r=0.5$, $\alpha=1$, $\zeta_{0}=0$, and monitor function $g(x,p)=\|\nabla U(x)\|$.
@@
We also perform more extensive runs to resolve the bias in canonical averages. The procedure is similar to the case of the 1D toymodel in Sec. 5.1. To obtain different $\langle\Delta t\rangle$ for SamAdams, we varied $\Delta\tau$, keeping everything else fixed, and then ran BAOAB at fixed stepsizes set to the obtained $\langle\Delta t\rangle$ values for comparison. We averaged over 200 independent trajectories, where the number of iterations scaled linearly with $\Delta\tau$ (we used $10^{8}$ iterations at $\Delta\tau=0.001$, and varied $\Delta\tau$ from 0.00025 to 0.06). The computational cost of each BAOAB simulation was the same as for the corresponding SamAdams run. Fig. 10 shows the bias against (mean) stepsize $\Delta t$ for the two temperatures and coordinates. While SamAdams and BAOAB show comparable errors for small stepsizes, BAOAB becomes unstable at $\Delta t=0.0025$. SamAdams can still be run at $\langle\Delta t\rangle=0.01$ without any decline in accuracy. If a modest decline in accuracy is acceptable, it can be run at substantially larger steps. In fact, while the BAOAB curves in Fig. 10 stop at the last stable stepsize, SamAdams remained stable until the last examined $\Delta\tau$.